# Trust, Because You Can’t Verify: Privacy and Security Hurdles in Education Technology Acquisition Practices

### v1

@@ 5 Privacy Posture (RQ1):
HEIs have privacy policies that, at a high level, guide all institutional activities that may require the collection and processing of data from students and staffs, as well as the acquisition and use of EdTech from external vendors. While these policies mostly rely on federal privacy regulations, some HEIs also incorporate state laws or form data governance models with additional terms and conditions.
@@ 5.1 Inadequacy of Federal Level Policies
In the United States, information privacy is regulated by domain, for example, health information is regulated by HIPAA and FERPA is the main federal regulation protecting education records [2]. Our interviews revealed that most of the institutional policies are largely governed by federal guidelines, such as FERPA and HIPAA. However, most of our participants, $70\%(n=9)$, commented that these federal regulations lack adequate protection for sensitive data. For instance, P4 mentioned the need for a comprehensive, rather than sector-specific, privacy law:
@@
These existing laws also do not impose heavy penalties on the vendors for violations, as P5 underscored that often vendors may only get a “little reputational hit” in case of data misuse and added that:
@@
In response to these challenges, participants emphasized the necessity for advocacy efforts aimed at addressing deficiencies in federal data privacy legislation.
Takeaway: Inadequate Federal Level Policies Most HEIs rely on FERPA and HIPAA, meanwhile, vendors may often get away with only a “little reputational hit” in case of data breach or misuse. Participants advocate to address these deficiencies in federal data privacy regulations.
@@ 5.2 Low adoption of state legislation
This approach significantly helped enhance confidence in privacy measures during EdTech purchases for P11’s HEI.
@@
5.3 Dealing with intra- and inter-national policies.
Most HEIs have students, faculties, and staff from different parts of the world. Some HEIs discussed the challenges of navigating diverse state and international laws, particularly those with campuses spanning multiple states or enrolling international students. These institutions will “follow US law first” (P5), striving to comply with regulations from other countries such as the European Union’s General Data Protection Regulation (GDPR) [38], which imposes huge fines for non-compliance[39]. Although P5’s institution does not have campuses outside of the US, they discussed challenges towards adopting and prioritizing international laws:
@@
P4, whose institution offers online courses for international students, added “[HEIs] have to navigate a pretty quickly changing privacy space”. To proactively maintain compliance with the rapidly changing (international) privacy landscape, P4’s institution adopted the NIST privacy framework [40], invested in automation and tools like ServiceNow [^4] A cloud platform used for efficient collaboration and workflow across campus, https://www.servicenow.com/solutions/industry/education.html, and prioritized privacy protection in their data governance model. We discuss more about governance models in § 5.4.
@@ 5.4 Institutional data governance model
A privacy-focused data governance model can help HEIs proactively manage issues by controlling access and curbing abusive practices, and two institutes we interviewed are already doing that. P4 discussed the formation and workings of their governance model:
@@
“The provost’s office has a governance group that meets monthly [to discuss privacy] and that helps [them] form standards and policy and creates a community of practitioners… [They are] sending folks to privacy foundation specifically to help us build governance models. And, then, we spent last year building out a data and privacy governance model for the campus and brought in a really world-renowned expert to help us find how we were going to do data and privacy governance across the campus. A unique model because most people do data governance [only] and they don’t have privacy as a subset and we insisted that privacy was more [peer] to the data governments.”
@@
While it was unclear how “business need” is defined, and if such needs triumph over data subjects’ privacy needs as noted in [21, 22], in some institutions, internal policies from the governance body often dominate the contracts during EdTech acquisition. One of the participants (P8) shared:
@@
Only some HEIs focused on the development of governance models; not all HEIs have the capabilities to achieve such a standard due to factors such as funding and personnel.
@@ 6 Acquisition Phase (RQ2):
This section details EdTech acquisition and how security and privacy issues influence the procurement process. Several participants explained different stages of the acquisition process, from initial evaluations to full integration within the university’s systems, while P6 summarized them as follows:
@@ 6.1 Factors dominating the acquisition process
When asked what factors are considered during acquisition, participants emphasized the importance of tool features, robustness, reliability, and the longevity of a vendor and its business model (P7). For most HEIs, budget is a top priority (nine participants), as highlighted by P13:
@@
“The biggest challenge I have is budgets to be honest. You know, they want the top of the line, but they of course didn’t want to spend [the money] on it.” As expressed by P4, procurement processes involve various discussions on risk management:
{quoting}
“Procurement has a variety of different thresholds by which we go through different kinds of processes for procurement. Procurement will also be separate from IT risks. Although, we will have risk discussions that might include our subject counsel and an elevator procurement discussion and a lot of these cases. If it becomes something expensive we actually go out to an RFP (Request for Proposal) and their RFP processes then to support it.”
During the process, HEIs take a look at vendor’s capabilities. HEIs hold vendors who “enhance the students’ experience” (P1) and “[keeps] the learning in mind” (P8) as a top selling point. Following that is “how stable the product is” (P10). Security as a selling point is not fixated on until after the initial process of acquisition has begun.
On the other side, vendors often prioritize sales over security, as one of our participants (P5) noted:
@@
This indicates security is often not a priority during the initial stages of acquisition.
Takeaway: Factors impacting EdTech acquisition HEIs prioritize tool features and vendor capabilities and longevity, but face budget constraints. Security is a secondary concern—vendors prioritize sales over security.
@@ 6.2 Contractual Safeguards and Negotiations
Contracts serve as crucial tools for HEIs to regulate data collection and usage with vendors. HEIs critically examine vendor contracts, aiming to negotiate favorable data-use and liability terms and create long-term leverage. Thus, these negotiations are often time-demanding.
Data Ownership. HEIs prioritize maintaining the ownership of the data that EdTechs collect. According to P2, “it’s always university data” and P5 noted that “It’s pretty rare that we would allow vendors to claim ownership of it” and added that their “legal department pushes back if vendors claim ownership.” P3 shared a case where a vendor discontinued using student data after “concerns raised in the higher-ed space around privacy” and added that “we’re protected via contracts.”
Vendors, however, can license institutional data for previously agreed-upon uses, where they are responsible to for its secure handling. As P9 said, “it all comes down to the agreement.” P4 discussed how data sharing between vendors was handled using a “data transfer agreement or data processing agreement” laid out within the contract. P6 underscores the importance of contractual clarity in licensing data, cautioning unfavorable attempts of vendors claiming data ownership:
{quoting}
“So part of it is as most tools in order to work need access to and license of the content you are putting into them. But at the same time, we want to retain ownership of that content. We will grant you a license to use it in the scope of this agreement but that does not give you ownership of that content and that’s where sometimes there is a sloppy agreement that a vendor will send over that’s trying to take ownership of our IP [intellectual property] and… that’s just done on starters.”
Thus, the contractual language is a pillar stone for HEIs and the only safeguard against vendor issues.
Security, privacy and liabilities in contracts. Participants (P2, P3, P6, and P11) emphasized the importance of vendors agreeing to institutional terms, especially regarding security and privacy. P6 highlighted the process of negotiating terms to ensure compliance with institutional policies:
{quoting}
“Traditionally, we asked the vendor to agree to [the universities] terms and conditions… part of those terms and conditions includes recognition of security and privacy. And to assure that not only are they meeting them in the kind of short term, but that they continue to meet them, and that is a contingent part of our agreement going ahead”
All our participants’ institutions ensured that contracts clearly defined data ownership and the scope of data use. Often, contracts outline “confidentiality and privacy” to restrict vendors from “sell[ing] [institutional] data”. One of the participants (P1) states:
@@
Lastly, several participants indicated that HEIs struggle to ensure vendors comply with the contracts due to limited insight into vendor operations. P11 expressed how their office of general counsel was working towards finding out how to “get [checks on vendors] into the contracts, how do we enforce it and how do we check [that they are doing it]?”
@@
“There’s concern in the larger community with AI. So when we enable AI tools which students want to be able to use, faculty want to use, we want to do research on them. How much of that is actually going back and being used? So is it used to train the model? Where is it stored? What do you do with it? Do you share it with anybody? That’s kind of the challenge that I have right now with AI”
@@ 6.3 Pre-acquisition Security Assessments
Risk-driven security assessment. Since security assessment is time-intensive and often involves iterative exchanges, several participants (P4, P5, P6, P7, P8, P9, and P13) highlighted the importance of determining the risk profile of each tool early in the acquisition process. P6 emphasized that their institution’s first step is “to determine the risk profile of a tool” to tailor the intensity of the subsequent security assessment accordingly. P3 expressed that with smaller vendors “there’s a lot of growth and a lot of immaturity in terms of how those organizations deal with securing privacy.”
P2 explained how different vendors will have different assessments based on the information they are handling:
@@
“It’s different levels of rigor. A relatively small product with no significant data, we’re really focused on assessing security and whether they’re going to protect the integration of the service itself and make sure that it’s not a pivot place for places that hold data. Then there’s a more aggressive one, like we’re going to need to see your plans and get your outputs, make sure you send them along with your questionnaire report. If you’re filling out one of the big assessment reports, send those along so we can keep those on record. ”
Based on the risk profile assessment, HEIs ask vendors to provide audit reports (from either vendors’ own assessments or third-party audits like SOC2) or fill up questionnaires like HECVAT. As P2 stated:
{quoting}
“Most of what we do is ask for their outputs from their own assessments or an independent party’s assessment of their operation to verify that someone beyond them have said that they’re doing what they’re supposed to.”
@@
Despite these limitations, P12 acknowledged the utility of HECVAT assessments for a baseline level of trust:
{quoting}
“it [HECVAT] is useful to gain some trust in that they answered it, but again that’s why I like to look at the SOC 2 and the [penetration] test. Plus, if they answered something on there and then you know something happened. You sort of have that attestation from them that they were doing it”
P2, P3, and P8 similarly take extra measures of penetration tests to make sure that “they’re [EdTech] secure” (P3) and “we aren’t connecting any potential vulnerabilities in our systems” (P8). While this multifaceted approach to security assessment reflects institutions’ efforts to mitigate security risks, some EdTech may escape any assessments if they do not go through institutional procurement departments (e.g., being adopted through a less centralized process by the HEI), creating additional vulnerabilities, as P11 disclosed it:
@@ 6.4 Sub-vendors and associated concerns
HEIs grapple with additional challenges of visibility and the “unknown” due to the involvement of sub-vendors (vendors’ vendors). This expansion opens the EdTech ecosystem to a larger threat landscape—necessitating HEIs to adapt to and adjust their processes accordingly. With more actors playing a part, the risk of unauthorized access to an institution’s data increases drastically, eventually requiring the HEIs to not only consider the risk profile of their vendor but also its sub-vendors. Five of our participants (P2, P3, P5, P11, and P12) found sub-vendor breaches concerning as they might lead to an institution’s data leak.
@@
P2 shared that the proliferation of sub-vendors with inadequate resources exacerbates the risks of sub-vendor breaches.
Additional measures for handling sub-vendors. At the start of an agreement, some HEIs (P11 and P12) require vendors to disclose their vendors’ information and mandate audits for sub-vendors based on the initial risk assessment. Some institutions have dedicated teams working on ensuring that tools from (sub-) vendors that plug into their system are safe by implementing a version of the product on a test bed, P13 said how they did it:
{quoting}
“If I’m deploying something new I’ll do like a test bed before for at least a month normally. Out there, test it, you know, do all the scenarios that I can think of to break it”
@@ 6.5 Collaboration among HEIs for better leverage
The lack of a clear picture of the vendors’ security and privacy posture, exacerbated by reliance on questionnaire-based assessments and limited system access, motivated HEIs to develop collaboration groups, such as Educause [43], Big 10 Academic Alliance [44], or the University of California Information Security Council  [45]. These groups, comprising of HEI security officers (CISOs) and other leadership personnel, facilitate the exchange of notes on security practices and experiences working with different vendors. All our study participants indicated leveraging connections with other universities and groups to vet vendors for purchasing. P1 described their collaborative experience as “highly interactive and supportive.”
@@
“each university has more of a specific focus on some things, where’s the others don’t. Like a lot of ours is like research and medical… There’s kind of [a] different focus [at] each [university] and so we at least kind of have our own requirements on IT stuff. So yes, it’s different all over the place.”
{quoting}
@@ 7 Post-Acquisition (RQ3):
In this section, we discuss the challenges that institutions perceive and experience as they transition to the post-acquisition phase of integrating EdTech in their university networks.

### v2

@@ 5 Privacy Posture (RQ1):
HEIs have privacy policies that, at a high level, guide all institutional activities that may require the collection and processing of data from students and staff, as well as the acquisition and use of EdTech from external vendors. While these policies mostly rely on federal privacy regulations, some HEIs also incorporate state laws or form data governance models with additional terms and conditions.
@@ 5.1 Inadequacy of Federal Level Policies
Our interviews revealed that most institutional policies are governed largely by domain-specific federal guidelines, such as FERPA [17] and HIPAA [21] for education and health records, respectively. However, most of our participants ($n=9$), commented that these federal regulations lack adequate protection for sensitive data. For instance, P4 mentioned the need for a comprehensive, rather than sector-specific, privacy law:
@@
These existing laws also do not impose sufficiently large penalties on the vendors for violations, as P5 underscored that often vendors may only get a “little reputational hit” in case of data misuse and added that:
@@
In response to these challenges, participants emphasized the necessity for advocacy efforts aimed at addressing deficiencies in federal data privacy legislation. Some examples include HEIs coming together to establish standards (P2), provide inputs to the legislation (P3), and eventually, create policies that stand the test of time (P5).
Takeaway: Inadequate Federal Level Policies Most HEIs rely on FERPA and HIPAA, meanwhile, vendors may often get away with only a “little reputational hit” in case of data breach or misuse. Participants advocated addressing these deficiencies in federal data privacy regulations.
@@ 5.2 Low Adoption of State Legislation
P11’s HEI enhanced confidence in data privacy measures by enforcing state-provided five questions—among them, two on security and one on accessibility—during EdTech purchases and noted that “If you [vendor] don’t answer them yes, we just say we can’t do business with you. So the state helped us there, in a way.”
@@
5.3 Dealing with Intra- and Inter-national Policies
Several participants discussed the challenges of navigating diverse state and international laws, particularly those with campuses spanning multiple states or enrolling international students. These institutions will “follow US law first” (P5), striving to comply with regulations from other countries such as the European Union’s General Data Protection Regulation (GDPR) [40], which imposes huge fines for non-compliance[41]. Although P5’s institution does not have campuses outside of the US, they discussed challenges towards adopting and prioritizing international laws:
@@ 5.3 Dealing with Intra- and Inter-national Policies
P4, whose institution offers online courses for international students, added “[HEIs] have to navigate a pretty quickly changing privacy space”. To proactively maintain compliance with the rapidly changing (international) privacy landscape, P4’s institution adopted the NIST privacy framework [42], invested in automation and tools like ServiceNow [^4] A cloud platform used for efficient collaboration and workflow across campuses, https://www.servicenow.com/solutions/industry/education.html, and prioritized privacy protection in their data governance model. We discuss more about governance models in § 5.4.
@@ 5.4 Institutional Data Governance Model
Five participants (P3, P4, P8, P11, and P12) described how their HEIs have developed privacy-focused data governance models, sometimes referred to as General Counsel, to proactively manage EdTech acquisition-related issues. These models review vendor terms for alignment with HEI privacy standards, control access, and curb abusive practices. P4 discussed the workings of their governance model:
@@
“The provost’s office has a governance group that meets monthly [to discuss privacy]. That helps [them] form standards and policy and creates a community of practitioners. [They are] sending folks to Privacy Foundation [^5] The Privacy Foundation mentioned here is a research institution that holds seminars for policy-makers, scholars, and other members of the community to educate people on privacy and policy issues faced by HEIs [43]. specifically to help us build governance models. And then, we spent last year building out a data and privacy governance model for the campus and brought in a really world-renowned expert to help us find how we were going to do data and privacy governance across the campus. A unique model because most people do data governance [only] and they don’t have privacy as a subset and we insisted that privacy was more [peer] to the data governments.”
@@
While it was unclear how “business need” was defined, and if such needs triumph over data subjects’ privacy needs as noted in [19, 20]. In some institutions, internal policies from the governance body often dominate the contracts during EdTech acquisition. One of the participants (P8) shared:
@@ 6 Acquisition Phase (RQ2):
This section details EdTech acquisition and how security and privacy issues influence the procurement process. Several participants outlined the intricate processes involved in procuring EdTech solutions, with P6 summarizing them as follows:
@@
This quote reflects broader experiences shared by other participants. For instance, P1, P2, and P3 particularly discussed the necessity of negotiating terms to ensure the inclusion of required features and appropriate pricing.
@@ 6.1 Factors Dominating the Acquisition Process
When asked what factors are considered during acquisition, participants emphasized the importance of tool features, robustness, and reliability ($n=9$), as well as the longevity of a vendor and its business model (P3, P7, P8). For most HEIs, budget is a top priority ($n=9$), as highlighted by P13:
@@
“The biggest challenge I have is budgets to be honest. You know, they want the top of the line, but they of course didn’t want to spend [the money] on it.”
HEIs also evaluate vendor capabilities, prioritizing those that “enhance the students’ experience” (P1) and “[keep] the learning in mind” (P8), followed by “how stable the product is” (P10). Security considerations typically arise later in the acquisition process. On the other end, vendors often prioritize sales over security, as noted by P5:
@@
This indicates security is neither a priority by vendors nor by HEIs (at least during the initial stages of acquisition).
Takeaway: Factors impacting EdTech acquisition HEIs prioritize tool features, vendor capabilities, and longevity, but face budget constraints. Security is a secondary concern. Vendors prioritize sales over security.
@@ 6.2 Contractual Safeguards and Negotiations
Contracts serve as crucial tools for HEIs to regulate data collection and usage with vendors. HEIs critically examine vendor contracts, aiming to negotiate reasonable data-use and liability terms and create long-term leverage. These negotiations are often time-demanding.
Data ownership. HEIs prioritize maintaining the ownership of the data that EdTechs collect via contracts. According to P2, “it’s always university data” and P5 noted that their “legal department pushes back if vendors claim ownership [of HEI data].” P3 shared a case where a vendor discontinued using student data after “concerns raised in the higher-ed space around privacy” and added that “we’re protected via contracts.” As P9 said, “it all comes down to the agreement.”
All our participants’ institutions ensured that contracts clearly defined data ownership and the scope of data use. Often, contracts outline “confidentiality and privacy” to restrict vendors from “sell[ing] [institutional] data”. P1 stated:
@@
Vendors, however, can license institutional data for previously agreed-upon uses and are responsible for its secure handling. To manage data sharing with vendors, P4 mentioned the use of “data transfer agreements” or “data processing agreements.” P6 underscored the importance of contractual clarity in licensing data, cautioning against unfavorable attempts of vendors to claim data ownership through “sloppy agreements.” Thus, the contractual language is a pillar stone for HEIs and the only safeguard against vendor issues.
P3 stated that vendors who understand that “data that goes into their systems is the property of [university]” is important. Therefore, according to P3, establishing strong partnerships between HEIs and vendors can be particularly beneficial in the evolving landscape of EdTech. This approach ensures that vendors are not only aware of but also aligned with the needs and expectations of HEIs, facilitating more effective and responsive data management.
Security and privacy terms in contracts. Participants (P2, P3, P6, and P11) emphasized the importance of vendors agreeing to institutional terms, especially regarding security and privacy. P6 highlighted the process of negotiating terms to ensure compliance with institutional policies:
{quoting}
“Traditionally, we asked the vendor to agree to [the universities] terms and conditions… part of those terms and conditions includes recognition of security and privacy. And to assure that not only are they meeting them in short term, but that they continue to meet them, and that is a contingent part of our agreement going ahead.”
@@
Lastly, several participants indicated that HEIs struggle to ensure vendors comply with the contracts due to limited insight into vendor operations (§ 7.2). P11 expressed how their office of general counsel was working towards finding out how to “get [checks on vendors] into the contracts, how do we enforce it and how do we check [that they are doing it]?”
@@
“So when we enable AI tools which students want to use, faculty want to use, we want to do research on them. How much of that is actually going back and being used? So is it used to train the model? Where is it stored? What do you do with it? Do you share it with anybody? That’s kind of the challenge that I have right now with AI”
@@ 6.3 Pre-acquisition Security Assessments
Risk-driven security assessment. Since security assessments are time-intensive and often involve iterative exchanges, several participants highlighted the importance of determining the risk profile of each tool early in the acquisition process. P6 emphasized that their institution’s first step is “to determine the risk profile of a tool” to tailor the intensity of the subsequent security assessment accordingly. P2 explained how different vendors will have different assessments based on the information they are handling:
@@
“It’s different levels of rigor. A relatively small product with no significant data, we’re really focused on assessing security and whether they’re going to protect the integration of the service itself and make sure that it’s not a pivot place for places that hold data. Then there’s a more aggressive one, like we’re going to need to see your plans and get your [assessments] outputs, make sure you send them along with your questionnaire report. If you’re filling out one of the big assessment reports, send those along so we can keep those on record. ”
Based on the risk profile assessment, HEIs ask vendors to provide audit reports (from either vendors’ own assessments or third-party audits like SOC2) or fill up questionnaires like HECVAT. P2 asks vendors “for their outputs from their own assessments or an independent party’s assessment of their operation to verify that someone beyond them have said that they’re doing what they’re supposed to.”
@@
Despite these limitations, P12 acknowledged that “[HECVAT] is useful to gain some trust in that they answered it, but again that’s why I like to look at the SOC 2 and the [penetration] test.” P2, P3, and P8 similarly take extra measures of penetration tests to make sure that “they’re [EdTech] secure” (P3) and “we aren’t connecting any potential vulnerabilities in our systems” (P8).
While this multifaceted approach to security assessment reflects institutions’ efforts to mitigate security risks, some EdTech may escape any assessments if they do not go through institutional procurement departments (e.g., being adopted through a less centralized process by the HEI), creating additional vulnerabilities, as P11 disclosed it:
@@ 6.4 Sub-vendors and Associated Concerns
HEIs grapple with additional challenges of visibility and the “unknown” due to the involvement of sub-vendors (vendors’ vendors). This expansion opens the EdTech ecosystem to a larger threat landscape—necessitating HEIs to adapt and adjust their processes accordingly. With more actors playing a part, the risk of unauthorized access to an institution’s data increases drastically, eventually requiring the HEIs to not only consider the risk profile of their vendor but also its sub-vendors. Five of our participants (P2, P3, P5, P11, and P12) found sub-vendor breaches concerning as they might lead to an institution’s data leak.
@@
Additional measures for handling sub-vendors. At the start of an agreement, some HEIs require vendors to disclose their vendors’ information and mandate audits for sub-vendors based on the initial risk assessment. Some institutions, such as P13’s, have dedicated teams working on ensuring that tools from (sub-) vendors that plug into their system are safe by implementing a version of the product on a test bed.
@@ 6.5 Collaboration among HEIs for Better Leverage
The lack of a clear picture of the vendors’ security and privacy posture, exacerbated by reliance on questionnaire-based assessments and limited system access, motivated HEIs to develop collaboration groups, such as Educause [46], Big 10 Academic Alliance [47], or the University of California Information Security Council  [48]. These groups, comprising of HEI security officers (CISOs) and other leadership personnel, facilitate the exchange of notes on security practices and experiences working with different vendors.
{quoting}
@@
Information sharing after privacy incidents. P3 described how they use their collaboration with other universities to help inform them about privacy incidents and how to make vendors follow necessary obligations:
{quoting}
“We had an incident in the recent past [with data leakage] that we, let’s just say, leveraged the REN-ISAC that the community is sharing such that others were made aware of some of the issues of this third party. So not only do we leverage the contract and enforce upon the vendor their obligations to notify other institutions that may have been impacted and notify students that may have been impacted, but also leverage that community network that we have to help inform others so that they know the extent of what’s happened.”
@@
“Each university has more of a specific focus on some things, whereas the others don’t. Like a lot of ours is like research and medical… There’s kind of [a] different focus [at] each [university] and so we at least kind of have our own requirements on IT stuff. So yes, it’s different all over the place.”
@@ 7 Post-Acquisition (RQ3):
In this section, we discuss the challenges that institutions perceive and experience as they transition to the post-acquisition phase of integrating EdTech into their university networks.